# Multilingual Prompting for Improving LLM Generation Diversity

### v1

@@ 4.1 Experimental Setup
Prompting Comparisons, Baselines, and Performance Tests. To ensure a fair comparison across prompting strategies, we generate three LLM responses with each strategy (multicultural, multilingual, and the baselines and comparison methods below), and evaluate the diversity of the concatenated responses. The exact phrasing of all prompts is included in Appendix A.1.2 and Appendix A.2.1.
Baseline. Our baseline consists of prompting the model in English with queries from each of the datasets above, each time with a preamble stating that the LLM is a helpful assistant. We refer to this as monolingual prompting. In order to create some amount of diversity across the generations in this strategy, we rephrase the prompts into multiple distinct variants with changes in phrasing, syntax, and placement of clauses, as prior work has shown that models can be quite sensitive to these attributes Sclar et al. (2023). (See more details in Figure 2, Appendix A.1.2 and  A.2.1).
Comparisons. To assess the effectiveness of our approach, we compare our method against previously established diversity-enhancing techniques: (1) High-temperature sampling, using the monolingual strategy from above, but setting temperature = 1.3 Chung et al. (2023). (2) Requesting Diversity: We also compare with prompts that simply ask the model to be diverse, namely by adding “Please try to be as diverse as possible” to the monolingual prompt. For these two comparison methods, to increase diversity, we also evaluate the diversity over concatenated responses of three rephrased versions of the prompt. (3) Random Personas: Following prior work Wang et al. (2025), we create personas for the model prior to prompting. To separate persona prompting from multilingual prompting, these prompts do not encode cultural information, but rather professions and other personality traits. We use the same number of personas as languages and evaluate concatenated responses. (4) Step-by-step Recall Hayati et al. (2023): This prepends past answers to subsequent questions sequentially to ask the model to generate new answers after reflection on prior answers. To compare fairly, we generate query responses from three rounds of Step-by-Step Recall, and evaluate the concatenated responses.
@@
We include Step-by-Step Recall and Requesting Diversity for the demographic diversity tasks but not social norm tasks, as they do not work well with multiple choice outputs. Step-by-Step Recall asks the model to reveal its first answer and then generate a different one in the next round, forcing the model to change its mind, which contradicts the spirit of a single-choice multiple-choice task. Similarly, Requesting Diversity is designed to elicit a set of varied outputs, but in the social-norm setting the model must commit to exactly one label, so the notion of “being diverse” reduces to a single token and loses its intended effect.
@@
Performance Checks. To ensure that the LLMs are reasoning well when responding to the multiple-choice questions given from the modified social norms dataset Forbes et al. (2020), we perform a test with different multiple-choice responses based on Zellers et al. Zellers et al. (2019), where three our of four responses are logically nonsensical reasons for agreeing or disagreeing to the controversial statement, discussed further in Appendix A.8.3. More broadly, to verify that multilingual prompting does not compromise the factual accuracy of language models, we evaluate their performance on the Multilingual Grade School Math Benchmark (MGSM) Shi et al. (2022), which consists of mathematical reasoning tasks translated into multiple languages. Across all models, we observe that multilingual prompting maintains similar factual accuracy to monolingual prompting: GPT-4o-mini shows virtually no change; for GPT-4o and LLaMA-70B, there is a slight performance drop around 5%, but the overall competency of the model remains intact. More information is in Appendix A.6.
@@
Models. We conduct experiments over four models: GPT-4o, GPT-4o-mini Hurst et al. (2024), LLaMA 3.3 70B and LLaMA 3.1 8B(Grattafiori et al., 2024).
@@ 4.2 Results
Multilingual Prompting Boosts Diversity of LLM Responses. To evaluate whether and how multilingual and multicultural prompting promotes more opinion diversity across social norm-related questions, and demographic diversity in questions about individuals, we compare LLM responses across prompting strategies using the three metrics defined earlier: Reason Entropy, Agreement Entropy, and Demographic Entropies. Table 1 reports the mean normalized entropy scores for each model across the different prompting strategies. Strategies are grouped into baseline, comparison, and multilingual and multicultural (our) methods. Due to space constraints, we present the average results for nationality, ethnicity, and geographic region diversity. Full results, including graphs of table results for ease of interpretation, are in Appendix A.8.
Beyond outperforming comparison methods and baselines, multilingual prompting methods consistently outdo multicultural prompting methods, suggesting the added importance of language in reaching different regions of an LLM’s knowledge base. Interestingly, the added benefit of language vary depending on the level of added cultural cues in the prompts: language is especially helpful when there is less cultural information in the prompt. Basic multilingual prompting performs markedly better than basic multicultural, by a factor 2x on average for reasoning and agreement entropy (social norm) experiments and 1.1x on averaged demographic entropies. Meanwhile, with the exception of two outliers from LLaMA-8B, enhanced multilingual only outperforms enhanced multicultural by a factor of 1.09 on average for reasoning and agreement entropy (social norm) experiments and 1.04x on averaged demographic entropies. These results suggest that language and cultural cues are both important components of eliciting diverse responses, but that they are best together (i.e., enhanced multilingual performs the best). This may be surprising given prior work showing minimal impact of language in eliciting specific cultural perspectives Kwok et al. (2024), but aligns with prior work suggesting that LLMs have language-specific knowledge bases Aggarwal et al. (2025).
Thus, our results suggest that both linguistic variation and cultural cues in input prompts serve as a valuable signal for models to generate more inclusive and varied content, reflecting a broader range of perspectives and cultural attributes. Using these cultural and language cues are significantly better at eliciting diverse responses from the model than other diversity-enhancing methods. In the next section, we show that beyond mild gains in improving diversity, multilingual prompting performs better than multicultural prompting, as we see that multilingual prompting prevents hallucination about culturally-relevant information.
@@
Table 1: Normalized entropy across social norm (Reason, Agreement) and demographic representation (Demo Avg.). “Demo Avg.” stands for the demographic average between nationality, ethnicity, and region. ‘—’ indicates experiments not run, explained in Section 4.1. * indicate the statistically significant differences between our methods and the best performance in Diversity-Enhancing.
@@ 5 Language Helps Prevent Hallucination
We now demonstrate that language is an important component of multilingual prompting, as it leads to lower hallucination rates for non-English speaking cultures. In particular, we demonstrate that multicultural prompting with cultural cues but without including the relevant language, (i.e., “Chinese-speaking, born in Bejing” but the prompt is not in Chinese) can lead to higher hallucination rates on non-Western names in queries about individuals.
@@ 5.1 Experimental Setup
For this experiment, we test whether Chinese names generated in response to questions about individuals in various professions are hallucinated or not. To do so, we first randomly sample profession-name pairs generated by the Chinese component of the (basic and enhanced) multilingual and multicultural prompting strategies on the People Diversity Dataset, which asks about naming individuals from different professions. We specifically sample from the subset which were annotated as Chinese by the labeling LLM. We sample 105 pairs each for the basic multilingual, multicultural, enhanced multilingual and enhanced multicultural methods.
Then, to calculate the hallucination rate of generated names, we collect human annotations through Prolific. We say a name is hallucinated for a given profession query if a that name is not associated with a person in that profession through Google or Wikipedia search. Annotators are given a name and profession from the LLM generation. They are instructed to search the name on Google and Wikipedia, and report whether the name is likely a hallucination i.e., not associated with someone of that profession, or not. To ensure accuracy, each name is evaluated independently by three different annotators. Authors manually inspect inconsistent cases (details in Appendix A.7).
@@ 5.2 Results
Language Helps Prevent Hallucination. The evaluation reveals a notable difference between the hallucination rate of Chinese names generated from a prompt in Chinese, versus in English. The multilingual strategy (using Chinese prompts) achieves an validity rate of 92 out of 105 (87.6%), whereas the multicultural strategy (using English) attains a lower rate of 77 out of 105 (73.3%). This 14% absolute improvement suggests that using the relevant language to cue the model to provide responses about a given culture is an important component of generating factually correct diverse responses. Moreover, the enhanced multilingual strategy (using Chinese prompts) achieves an validity rate of 97 out of 105 (92.3%), whereas the enhanced multicultural strategy (using English) attains a lower rate of 85 out of 105 (81.0%). These results confirm a trend seen in prior work, which has shown that LLMs have different factual knowledge across different languages Aggarwal et al. (2025).
@@ 6.1 Experimental Setup
To ensure that our high- versus lower-resourced experiments remain directly comparable across the $k=3,4,5$ language settings, we standardize both the amount of data collected and the scale on which each diversity metric is reported. Details on how this is done are in Appendix A.3. To ensure that models performed sufficiently well on lower-resourced languages to include in this experiment, we extend our performance check from Section 4.1 to lower-resourced languages, as well as testing instruction following. Results are in Appendix A.8.3. GPT-4o and GPT-4o-mini perform well, and LLaMA-70B and 8B do not, so we do not include them.
@@ 6.2 Interaction Effects between Model Size and Resource Level
Further, our results reveal that diversity performance across low and high resource languages are model specific. For the larger GPT-4o model, high-resourced language combinations consistently yield higher diversity scores across all three metrics—Reason Entropy, Agreement Entropy, and Perspective Diversity. In contrast, for the smaller GPT-4o-mini model, lower-resourced language combinations outperform high-resource ones.
@@ 7 Conclusion
We introduce multilingual and multicultural prompting methods to enhance cultural diversity in LLM-generated responses. We show that they out-perform existing methods for this task. Moreover, we find that multilingual prompting is more effective than multicultural prompting, both for promoting diversity and for reducing model hallucination about culture-specific information—suggesting that language is an important component in eliciting more diverse responses.
Limitations
Finally, we discuss some limitations of our work. Broadly, enhancing diversity may not always be a good outcome— for example, we may not want a model to express diverse opinions on “it is okay to murder someone”. Establishing when is the right time to elicit diverse responses is out of scope for this work, but we look forward to exploring in future work.
Another limitation of our work is that we only explore concatenation as an aggregation strategy— for tasks which require succinct answers, summarization and random selection from a distribution would be a better choice. While we believe random selection would give identical results in aggregate, fully exploring how to synthesize the diverse perspectives and pieces of information generated through multilingual prompting requires more study, which we look forward to in future work.
Further, language translation represents another potential source of limitation. While the authors possess fluency in English, Chinese, and Japanese, translations involving other languages were conducted using GPT models (GPT-4o). Existing evaluations and our empirical observations commonly suggest that GPT achieves near-human performance in translation tasks; however, subtle semantic or cultural nuances may not be fully captured in some instances.
Additionally, to mitigate reproducibility concerns and reinforce the transparency of our findings, we have included the complete set of prompts and additional experimental outputs in the appendix. The supplementary materials are intended to facilitate the verification of our results and support the trustworthiness of our conclusions.
@@ A.1.1 Generating Choices
We show two representative outputs in Table 2, each including two agreeing and two disagreeing perspectives.
@@ A.3 Metric Normalization for Social Norm Experiemnt
so that $r$ categories receive $q+1$ answers and the remaining $m-r$ categories receive $q$ answers. The corresponding empirical probabilities are
@@
(We adopt the convention $00\log 0:=$ 0 whenever a probability is zero.)
@@
    $k=5$ $:~{}H_{\max}=-\!\Bigl{(}\tfrac{2}{5}\log\tfrac{2}{5}+3\,\tfrac{1}{5}\log\tfrac{1}{5}\Bigr{)}\approx 1.332.$
@@
    $k=3$ $:~{}H_{\max}=-\!\bigl{(}\tfrac{1}{3}\log\tfrac{1}{3}+\tfrac{2}{3}\log\tfrac{2}{3}\bigr{)}\approx 0.637,$
    $k=4$ $:~{}H_{\max}=\log 2\approx 0.693,$
    $k=5$ $:~{}H_{\max}=-\!\bigl{(}\tfrac{2}{5}\log\tfrac{2}{5}+\tfrac{3}{5}\log\tfrac{3}{5}\bigr{)}\approx 0.673.$
  • Perspective Diversity (a.k.a. Perspective Entropy). For each statement we embed the four choices =$\mathcal{E}=\{\mathbf{e}_{A},\mathbf{e}_{B},\mathbf{e}_{C},\mathbf{e}_{D}\}$ using Sentence-BERT. With $k$ languages ($k\!\in\!\{3,4,5\}$), consider every size-$k$ subset $S\subseteq\mathcal{E}$. For any subset $S=\{i_{1},\dots,i_{k}\}$ we define its mean pairwise dissimilarity
@@
    i.e. the largest dissimilarity obtainable from any size-$k$ subset.
@@
    Averaging $\widetilde{H}_{\text{Persp}}(q,k)$ over all statements places the metric on the common $[0,1]$ scale: $1$ indicates the greatest possible diversity, 00 indicates none.

### v2

@@ 4.2 Baselines and Performance Tests
To ensure a fair comparison across prompting strategies, we generate three LLM responses with each strategy (multicultural, multilingual, and the baselines and comparison methods below), and evaluate the diversity of the concatenated responses. The exact phrasing of all prompts is included in Appendix A.1.2 and Appendix A.2.1.
Baseline. Our baseline consists of prompting the model $n$ times in English only. To ensure comparability with multilingual methods, we adopt the same sampling strategy: each query is preceded by a short preamble (“You are a helpful assistant”) and rephrased into multiple lexical variants with high syntactic similarity, as prior work has shown that LLMs are sensitive to phrasing Sclar et al. (2023). Then, the $n$ outputs are concatenated and used to compute diversity metrics. We refer to this as monolingual prompting. By comparing such monolingual and multilingual prompting, we can attribute observed entropy gains to cross-lingual variation rather than the phrasing sensitivity of LLMs in generation. (See Figure 2 and Appendix A.1.2 and A.2.1 for prompt details.)
Comparisons.
To assess the effectiveness of our approach, we compare our method against previously established diversity-enhancing techniques:
  • High-temperature sampling, using the monolingual strategy from above, but setting temperature = 1.3 Chung et al. (2023).
  • Requesting Diversity: We also compare with prompts that simply ask the model to be diverse, namely by adding “Please try to be as diverse as possible” to the monolingual prompt. For these two comparison methods, to increase diversity, we also evaluate the diversity over concatenated responses of three rephrased versions of the prompt.
  • Random Personas: Following prior work Wang et al. (2025), we create personas for the model prior to prompting. To separate persona prompting from multilingual prompting, these prompts do not encode cultural information, but rather professions and other personality traits. We use the same number of personas as languages and evaluate concatenated responses.
  • Step-by-step Recall Hayati et al. (2023): This prepends past answers to subsequent questions sequentially to ask the model to generate new answers after reflection on prior answers. To compare fairly, we generate query responses from three rounds of Step-by-Step Recall, and evaluate the concatenated responses.
@@
We include Step‑by‑Step Recall and Requesting Diversity for the demographic diversity tasks, but not social norm tasks, as they do not work well with multiple-choice outputs. Step‑by‑Step Recall asks the model to reveal its first answer and then generate a different one in the next round, forcing the model to change its mind, which contradicts the spirit of a single‑choice multiple‑choice task. Similarly, Requesting Diversity is designed to elicit a set of varied outputs, but in the social‑norm setting, the model must commit to exactly one label, so the notion of “being diverse” reduces to a single token and loses its intended effect.
Performance Checks.
@@
To ensure that the LLMs are reasoning well when responding to the multiple-choice questions given from the modified social norms dataset Forbes et al. (2020), we perform a test with different multiple-choice responses based on Zellers et al. Zellers et al. (2019), where three out of four responses are logically nonsensical reasons for agreeing or disagreeing to the controversial statement, discussed further in Appendix A.8.3. More broadly, to verify that multilingual prompting does not compromise the factual accuracy of language models, we evaluate their performance on the Multilingual Grade School Math Benchmark (MGSM) Shi et al. (2022), which consists of mathematical reasoning tasks translated into multiple languages. Across all models, we observe that multilingual prompting maintains similar factual accuracy to monolingual prompting: GPT-4o-mini shows virtually no change; for GPT-4o and LLaMA-70B, there is a slight performance drop around 5%, but the overall competency of the model remains intact. More information is in Appendix A.6.
Models.
@@
We conduct experiments over four mainstream models: GPT-4o, GPT-4o-mini Hurst et al. (2024), LLaMA 3.3 70B and LLaMA 3.1 8B (Grattafiori et al., 2024).
@@ 4.3 Results
Table 1: Normalized entropy across social norm (Reason, Agreement) and demographic representation (Demo Avg.). “Demo Avg.” stands for the demographic average between nationality, ethnicity, and region. ‘—’ indicates experiments not run, explained in Section 4.1. * indicates the statistically significant differences between our methods and the best performance in diversity-enhancing comparisons.
Multilingual Prompting Boosts Diversity of LLM Responses.
To evaluate whether and how multilingual and multicultural prompting promotes more opinion diversity across social norm-related questions, and demographic diversity in questions about individuals, we compare LLM responses across prompting strategies using the three metrics defined earlier: Reason Entropy, Agreement Entropy, and Demographic Entropies. Table 1 reports the mean normalized entropy scores for each model across the different prompting strategies. Strategies are grouped into baseline, comparison, and multilingual and multicultural (our) methods. Due to space constraints, we present the average results for nationality, ethnicity, and geographic region diversity. Full results, including graphs of table results for ease of interpretation, are in Appendix A.8.
Beyond outperforming comparison methods and baselines, multilingual prompting methods consistently outdo multicultural prompting methods, suggesting the added importance of language in reaching different regions of an LLM’s knowledge base. Interestingly, the added benefit of language varies depending on the level of added cultural cues in the prompts: language is especially helpful when there is less cultural information in the prompt. Basic multilingual prompting performs markedly better than basic multicultural, by a factor 2x on average for reasoning and agreement entropy (social norm) experiments and 1.1x on average for demographic entropies. Meanwhile, with the exception of two outliers from LLaMA-8B, enhanced multilingual only outperforms enhanced multicultural by a factor of 1.09 on average for reasoning and agreement entropy (social norm) experiments and 1.04x on average demographic entropies.
Another interesting phenomenon we observe is that occasionally, basic multilingual prompting yields higher demographic entropy than enhanced multilingual prompting. We suggest that this is the result of a narrowing effect where stronger cultural cues for a specific cultural background in enhanced multilingual prompting can lead the model to elicit responses centered around that particular region or culture, sometimes reducing overall diversity. By contrast, the basic multilingual prompts are more likely to be a mix of multiple cultures or regions, thus yielding higher entropy. We see support for this idea when analyzing the names generated across the two kinds of prompts. For example, when comparing enhanced and basic multilingual prompting in Chinese, GPT-4o produces 547 Chinese names with the enhanced multilingual prompt versus 427 with the basic multilingual prompt out of 1050 names generated. Similarly, GPT-4o-mini produced 597 versus 369 Chinese names, and LLaMA-8B produced 728 versus 293 Chinese names in enhanced multilingual versus basic multilingual, respectively. This behavior may suggest that basic multilingual prompting can be preferable for achieving broad demographic coverage, whereas enhanced multilingual prompting may be more suitable when culturally specific responses are desired.
Taken together, our results suggest that both linguistic variation and cultural cues in prompts serve as valuable signals for models to generate more inclusive and varied content, reflecting a broader range of perspectives and cultural attributes. Further, our results show that using these cultural and language cues together is a more effective strategy for eliciting diverse responses from the model than other diversity-enhancing methods. These results may be surprising given prior work showing minimal impact of language in eliciting specific cultural perspectives Kwok et al. (2024), but align with prior work suggesting that LLMs have language-specific knowledge bases Aggarwal et al. (2025). In the next section, we show that beyond mild gains in improving diversity, multilingual prompting performs better than multicultural prompting, as we see that multilingual prompting prevents hallucination about culturally-relevant information.
@@ 5 Language Helps Prevent Hallucination
Figure 4: Error rates of Chinese names generated under two prompting strategies. Using multilingual prompts in Chinese yields a lower error rate compared to multicultural prompts (cultural cues but without including the relevant language) in English, demonstrating that prompting in the relevant language reduces hallucination.
We now demonstrate that language is an important component of multilingual prompting, as it can lead to lower hallucination rates for recalling information related to cultures and nationalities where English is not a main language. In particular, we demonstrate that multicultural prompting with cultural cues but without including the relevant language, (i.e., “Chinese-speaking, born in Bejing” but the prompt is not in Chinese) can lead to higher hallucination rates on non-Western names in queries about individuals. For example, in the multicultural setting, when asked, “Who are some circus performers that you admire?”, the model responded with “Zhang Yimou.” However, Zhang Yimou is not a circus performer but rather a renowned Chinese film director. Such errors highlight how excluding the relevant language in prompts can increase hallucinations.
@@ 5.1 Experimental Setup
For this experiment, we test hallucination rates on Chinese names generated in response to questions about individuals in various professions. To do so, we first randomly sample profession-name pairs generated by the Chinese prompt component of the (basic and enhanced) multilingual and multicultural prompting strategies on the People Diversity Dataset, which asks about naming individuals from different professions. Within the profession-name pairs generated by the Chinese modified prompt, we specifically sample profession-name pairs that were annotated as Chinese by the labeling LLM. We sample 105 pairs each for the basic multilingual, multicultural, enhanced multilingual, and enhanced multicultural methods.
Then, to calculate the hallucination rate of generated names, we collected human annotations through Prolific. We classify a name as hallucinated for a given profession query if that name is not associated with a person in that profession through a Google or Wikipedia search. Annotators were given a name and profession from the LLM generation. They are instructed to search the provided name on Google and Wikipedia, and report whether the name is likely a hallucination i.e., not associated with someone of that profession, or not. To ensure annotation accuracy from Prolific annotators, each name is evaluated independently by three different annotators. Authors manually inspect inconsistent cases (details in Appendix A.7).
@@ 5.2 Results
Language Helps Prevent Hallucination. The evaluation reveals a notable difference between the hallucination rate of Chinese names generated from a prompt in Chinese, versus in English. The multilingual strategy (using prompts in Chinese with Chinese cultural cues) achieves an error rate of 13 out of 105 (12.4%), whereas the multicultural strategy (using prompts in English with Chinese cultural cues) attains a higher rate of 28 out of 105 (26.7%). This 14% difference suggests that using the relevant language to signal the model to provide responses about a given culture is an important component of generating diverse responses that are also factually correct. Moreover, the enhanced multilingual strategy (using prompts in Chinese with more elaborate Chinese cultural cues) achieves an error rate of 8 out of 105 (7.7%), whereas the enhanced multicultural strategy (using prompts in English with more elaborate Chinese cultural cues) attains a higher rate of 20 out of 105 (19.0%). These results confirm a trend seen in prior work, which has shown that LLMs have different factual knowledge across different languages Aggarwal et al. (2025).
@@ 6.1 Experimental Setup
To ensure that our high‑ versus lower‑resourced experiments remain directly comparable across the $k=3,4,5$ language settings, we standardize both the amount of data collected and the scale on which each diversity metric is reported. Details on how this is done are in Appendix A.3. To ensure that models performed sufficiently well on lower-resourced languages to include in this experiment, we extend our performance check from Section 4.1 to lower-resourced languages, as well as testing instruction following. Results are in Appendix A.8.3. GPT-4o and GPT-4o-mini perform well, and LLaMA-70B and 8B do not, so we do not include them.
@@ 6.2 Interaction Effects between Model Size and Resource Level
Further, our results reveal that diversity performance across low and high resource languages is model-specific. For the larger GPT-4o model, high-resourced language combinations consistently yield higher diversity scores across all three metrics—Reason Entropy, Agreement Entropy, and Perspective Diversity. In contrast, for the smaller GPT-4o-mini model, lower-resourced language combinations outperform high-resource ones.
@@ 7 Conclusion
We introduce multilingual and multicultural prompting methods to enhance cultural diversity in LLM-generated responses. We show that they outperform existing methods for this task. Moreover, we find that multilingual prompting—where the language matches the cultural cues added to each modified version of the original LLM query—is more effective than multicultural prompting—which simply provides cultural cues for various cultures but maintains all modified prompts in English— both for promoting diversity, and for reducing model hallucination about culture-specific information. This suggests that language is an important component in eliciting more diverse responses. We hope that our method can be an easy, accessible way to increase LLM generation diversity for relevant tasks.
8 Limitations
Finally, we discuss some limitations of our work. Broadly, enhancing diversity may not always be a good outcome. Establishing when is the right time to elicit diverse responses is out of scope for this work, but we look forward to exploring in future work.
Another limitation of our work is that we only explore concatenation as an aggregation strategy— for tasks which require succinct answers, other aggregation strategies, such as summarizing all answers from multilingual prompt components, or random selection from a distribution of generated responses to different prompt components would be a better choice. While we believe random selection would give identical results in aggregate, fully exploring how to synthesize the diverse perspectives and pieces of information generated through multilingual prompting requires more study, which we look forward to in future work.
Further, language translation represents another potential limitation. While the authors possess fluency in English, Chinese, and Japanese, translations involving other languages were conducted using GPT models (GPT-4o). Existing evaluations and our empirical observations commonly suggest that GPT achieves near-human performance in translation tasks; however, subtle semantic or cultural nuances may not be fully captured in some instances.
Additionally, to ensure reproducibility and reinforce the transparency of our findings, we have included the complete set of prompts and additional experimental outputs in the appendix. The supplementary materials are intended to facilitate the verification of our results and support the trustworthiness of our conclusions.
9 Acknowledgments
We would like to thank Chuhan Ku for the help on data annotation, Falaah Arif Khan for valuable discussions about the evaluation, and Eunsol Choi for the conceptual design. We would also thank the constructive comments from reviewers. This material is based upon work supported by the National Science Foundation (NSF) under Grant No. IIS-2239862.
@@ Appendix A Appendix
Appendix Roadmap
This appendix provides supplementary details supporting our main paper. Appendix A.1 and  A.2 describes the exact prompts used in the social norms and cultural representation experiments. Appendix A.2.2 outlines our annotation procedure for demographic attributes of generated names. Appendix A.3 and  A.4 explains the normalization and calculation of our diversity metrics. Appendix A.5 presents full experimental results (including tables and figures) that extend the main text. Appendix A.8.3 and  A.9 details our performance and validity checks. Appendix  A.8 provide additioanl result and  A.10 discuss the computational cost of different methods.
@@ A.1.1 Generating Choices
We show two representative outputs in Table 2, each including two agreeing and two disagreeing perspectives. A potential avenue for future research in opinion generation involves the incorporation of Retrieval-Augmented Generation (RAG) techniques, which may enhance the grounding of generated opinions in identifiable sociocultural frameworks (e.g., Confucian ethics, modern family structures).
@@ A.3 Metric Normalization for Social Norm Experiemnt
so that $r$ categories receive $q+1$ answers and the remaining mrm-r categories receive $q$ answers. The corresponding empirical probabilities are
@@
(We adopt the convention $0\log 0:=0$ whenever a probability is zero.)
@@
[…]
@@
[…]
  • Perspective Diversity (a.k.a. Perspective Entropy). For each statement we embed the four choices $\mathcal{E}=\{\mathbf{e}_{A},\mathbf{e}_{B},\mathbf{e}_{C},\mathbf{e}_{D}\}$ using Sentence‑BERT. With $k$ languages ($k\!\in\!\{3,4,5\}$), consider every size‑$k$ subset $S\subseteq\mathcal{E}$. For any subset $S=\{i_{1},\dots,i_{k}\}$ we define its mean pairwise dissimilarity
@@
    i.e. the largest dissimilarity obtainable from any size‑$k$ subset.
@@
    Averaging $\widetilde{H}_{\text{Persp}}(q,k)$ over all statements places the metric on the common $[0,1]$ scale: $1$ indicates the greatest possible diversity, 0 indicates none.